Use algebra to find the point at which the line k(x) = -1/2 x + 25/8 intersects the line h(x) = 5/4 x + 155/16.
We're solving 2x2 systems using graphing but I don't recognize the k(x) and the h(x).

Answer by Alan3354(69443)

Use algebra to find the point at which the line k(x) = -1/2 x + 25/8 intersects the line h(x) = 5/4 x + 155/16.
We're solving 2x2 systems using graphing but I don't recognize the k(x) and the h(x).
-------------
k(x) and h(x) are functions of x.
To graph them, replace them with y:
y = -x/2 + 25/8
y = 5x/4 = 155/16
Same thing.
-----
Since they both equal y (or the function h(x) & k(x) are equal at the intersection),
-x/2 + 25/8 = 5x/4 + 155/16
Step 1, eliminate fractions - multiply by 16
-8x + 50 = 20x + 155
-105 = 28x
x = -15/4
-------
Sub for x in any function or eqn to find y, and h(x) & k(x)